OpenStudy (anonymous):

please help mee!!!! a plane is descending towards a runway in New York. The plane is flying an altitude of 8 miles and is 175 from the runway, as measured on the ground. What angle will the plane's path make with the runway? Round to your nearest hundredth

OpenStudy (joannablackwelder):

The altitude could be viewed as the "y" coordinate of a right triangle. And the distance from the runway the "x" coordinate. Thus the angle made by the plane's path with the runway would be opposite from the altitude and adjacent to the distance from the runway.

OpenStudy (joannablackwelder):

tanθ=opp/adj so θ=tan^-1(opp/adj)

OpenStudy (anonymous):

ohk thank you

OpenStudy (joannablackwelder):

No worries :)

OpenStudy (anonymous):

\[\frac{180 N\left[\sin ^{-1}\left(\frac{8}{175}\right)\right]}{\pi }=2.62 {}^{\circ} \]